# Group-Control Motion Planning Framework for Microrobot Swarms in a Global Field

### v1

@@ Abstract
This paper investigates how group-control can be effectively used for motion planning for microrobot swarms in a global field. We prove that Small-Time Local Controllability (STLC) in robot positions is achievable through group-control, with the minimum number of groups required for STLC being $\log_{2}(n+2)+1$ for $n$ robots. We then discuss the complexity trade-offs between control and motion planning. We show how motion planning can be simplified if appropriate primitives can be achieved through more complex control actions. We identify motion planning problems that balance the number of robot groups and motion primitives with planning complexity. Various instantiations of these motion planning problems are explored, with simulations to demonstrate the effectiveness of group-control.
@@
Micro Robots Global Field Control and Motion Planning.
@@ 1 Introduction
Microrobots have gained significant attention for their potential in medical, environmental, and industrial applications. Effective control mechanisms are crucial for enabling diverse functionalities. Various control methods have been developed to address the unique challenges faced by microrobots, including magnetic control [22, 23, 16], optical control [9, 2, 15, 19], and acoustic control [10, 1]. These systems are controlled by global fields, and as a result they are massively underactuated.
@@
Currently, parallel control of multiple microrobots in a global field is based on the design of robots with distinct physical characteristics, which results in different responses of each robot to the same control signal. The Global Control Selective Response (GCSR) paradigm introduced in [7, 8] relies on fabrication to make each robot respond appropriately to the control signal. Turning-rate Selective Control (TSC), introduced in [20], is a variation of GCSR and differentiates robots by explicitly designing them with different turning rates. These methods face challenges in scaling to larger swarms due to manufacturing complexities. Ensemble Control (EC), proposed by [3], leverages differences in linear velocity and turning rate parameters among robots stemming from randomness in fabrication to control the position of individual robots. Unfortunately, the ability to control individual robots is inherently limited by noise so this approach also scales poorly.
@@
This paper focuses on stress-engineered MEMS microrobots (MicroStressBot), originally developed in [6]. The swarm is powered by a uniform electrostatic field generated by a substrate, which means that all robots are controlled by a single global signal. Their size limits onboard control logic and power storage, making it essential to simultaneously coordinate large groups of microrobots for applications like micro-assembly or drug delivery. Our work explores the concept of on-board multi-stage Physical Finite-State Machines (PFSMs) introduced in [21]. PFSMs allow robots to be individually addressed and activated one by one. Our previous work [17] takes the idea of PFSMs further by introducing group-based control. In this paper, we investigate the trade-off between the complexity of motion planning and control within the group-control framework. In particular, we introduce several instances of control and motion planning problems and discuss their scalability.
@@
Motion planning for microrobots involves determining a sequence of movements to achieve specific goals while avoiding obstacles and collisions. Various methods exist, including graph-based methods like Dijkstra’s and $A^{*}$ search [13], sampling-based algorithms like Rapidly-exploring Random Trees (RRTs) [14] and optimization-based methods [11, 29]. Recent advances in machine learning, such as [24, 26, 27], leverage neural networks to speed up RRTs and control techniques for mobile robot navigation. Despite a wealth of solutions for motion planning problems, motion planning for microrobot swarms in a global field remains a challenging problem because the system is highly dimensional yet massively underactuated, having a single control signal.
@@
Fundamentally, when a microrobot swarm is controlled by a global control signal, the robots can be controlled individually only when each robot responds to the control signal differently. In this paper, we introduce the group-control framework to differentiate the motion of each microrobot in the swarm. We prove that Small Time Local Controllability (STLC) is achievable in robot positions given an appropriate group allocation, with the minimum number of groups required for STLC being $log_{2}(n+2)+1$ for $n$ robots. Additionally, we discuss the complexity trade-offs between control and motion planning. We identify motion planning problems that balance the number of robot groups and motion primitives with planning complexity. Various instantiations of these motion planning problems are explored, with simulations to demonstrate the scalability and effectiveness of the proposed algorithms.
@@ 2.1 MicroStressBot
Electrostatic stress-engineered MEMS microrobot (MicroStressBot) is a 120 $\mu$m $\times$ 60 $\mu$m $\times$ 10 $\mu$m mobile microrobot platform introduced in [6]. A MicroStressBot has two actuated internal degrees of freedom (DOF): an untethered scratch-drive actuator (USDA) that provides forward motion, and a steering-arm actuator that determines whether the robot moves in a straight line or rotates. A single MicroStressBot can have its arm either raised or lowered, depending on the voltage applied across a substrate formed by an electrode array. When high enough voltage is applied to the substrate, the arm is pulled into contact with the substrate and the robot rotates around the contact point. In contrast, when the voltage is reduced below a threshold, the arm is raised, and the robot moves forward.
MicroStressBot control has been successfully implemented in [8]. It has been shown that if the pull-down and release voltages of robots are different, they can be independently controlled. However, it is difficult to consistently manufacture the robots so that they respond in the desired way. As a result, the approach scales poorly.
@@ 2.2 Physical FSM Robots
One solution to dramatically improve the scalability of the microrobot swarm control is to make the robots respond to a temporal sequence of (a small number of) voltage levels rather than to the voltage directly. Finite State Machines (FSMs) can accept a set of input sequences [25] (sequences of control signal levels). Previously, in [21], we proposed the on-board MEMS Physical FSM (PFSM) that, upon the acceptance of a unique control signal sequence, causes the arm of the MicroStressBot to be pulled up or down. PFSMs can be constructed from several basic modules that are combined together and thus fabricated efficiently. In this work, we build on this idea to propose group-control. In particular, we use the fact that several PFSM modules, each corresponding to one group, can be combined together so each robot can belong to several groups.
@@ 2.3 Group-Control
The core idea of group-control is that by equipping the robots with PFSMs, they can be assigned to (several) different groups. When the activation sequence corresponding to a group is sent to the swarm, all the robots that belong to the group are activated. In this paper, a MicroStressBot being activated corresponds to its arm being pulled up so it can move forward (translate). By assigning each robot to a unique subset of groups, we can differentiate between the robots and make them respond differently to a sequence of inputs that activate all the groups one after the other, each time moving the robots belonging to the group for some distance.
Table 1 shows how six robots can be assigned to different groups. In order to make the selection of groups unique to each robot, if we have $n$ robots, we need $m=\log_{2}(n+2)+1$ groups ($n+2$ rather than $n$ because each robot needs to belong to at least one group so it can move forward, and no robot can belong to all groups so it can rotate). We add one more group where none of the robots translate, they all rotate in place. If we have $m$ groups, this special group will be the group $G_{m}$.
As can be easily seen from Table 1, group assignment corresponds to assigning a unique bit pattern (subset of the groups) to each robot, where each bit corresponds to a group (labeled as $G_{i}$, $i=1,2,3$ in Table 1). Each robot then belongs to the groups where the corresponding bit equals $1$. For example, robot $R_{1}$ belongs to group $G_{3}$. Instead, robot $R_{3}$ belongs to the groups $G_{2}$ and $G_{3}$, etc. The group $G_{4}$ is the group with all 00’s, representing the additional group where all the robots are rotating. It is clear that this method guarantees that each robot has a distinct group allocation. Also, note that each group (apart from the added group like $G_{4}$, which contains no robots and corresponds to the case where all the robots rotate) can contain at most $\frac{2^{m-1}-2}{2}=2^{m-2}-1$ robots.
@@
The group allocation can be realized by the on-board PFSM. At each time step $k$, only one group of robots is activated, and the robots belonging to that group translate (move forward) while the remaining robots rotate. We call a sequence of group selections an activation sequence.
If $m$ is the number of groups and $n$ is the number of robots, we would need $k=O(\log_{2}m)=O(\log_{2}(\log_{2}(n+2)+1))$ bits to select a group through a PFSM [21]. Thus, PFSMs for group-based control can be significantly simpler than in the case when each robot needs to be selected individually.
@@ 3.1 Dynamics
To describe group-control mathematically, we start with a dynamic model of a MicroStressBot. The robot can move freely on a horizontal plane, so its configuration space is Euclidean group $SE(2)$. We will describe the state of the robot $i$ with a vector $[x_{i},y_{i},\theta_{i}]^{T}$, where $x_{i}$ and $y_{i}$ are the Cartesian coordinates of the pivot point of the robot and $\theta_{i}$ is the robot orientation. The equations of motion are given by:
$\frac{d}{dt}\begin{bmatrix}x_{i}\\ y_{i}\\ \theta_{i}\end{bmatrix}=a_{i}\cdot\begin{bmatrix}\cos\theta_{i}\\ \sin\theta_{i}\\ 0\end{bmatrix}\cdot u+(1-a_{i})\cdot\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}\cdot\omega_{i},$ (1)
where $a_{i}\in\{0,1\}$ is the switching input that determines whether the robot is moving forward or turning in place, while $u\in\mathbb{R}^{+}$ and $\omega\in\mathbb{R}^{+}$ are the rates of forward motion and rotation, respectively, where $\omega_{i}=u/r_{i}$ with $r_{i}$ being the turning radius of each robot. If the control inputs are piecewise constant over each epoch $\Delta T$, we obtain the following discrete-time model:
@@
Note that $u$ is a unilateral control input. In other words, MicroStressBot can not go backward or turn counterclockwise.
@@ 3.2 Switched and Embedded Systems
By introducing new variables $v_{i}(t)\in[0,1]$ that satisfy the constraint $\sum_{i=1}^{m}v_{i}(t)=1$, the switched system (3) can be converted to the embedded form [4, 28]:
@@
where $u_{i}$ is the control input for each vector field $f_{i}$. It can be shown [4] that the set of trajectories of the switched system (3) is a dense subset of the set of trajectories of the embedded system (4). Consequently, any trajectory in the embedded system can be approximated by the switched system to any desired accuracy, which means that the controllability of the embedded system is equivalent to the controllability of the switched system.
@@ 4 Controllability Analysis
In order for the robots to navigate among obstacles, the system, in general, needs to be Small-Time Locally Controllable (STLC). There are otherwise configurations in the obstacle-free configuration space $C_{free}$ that will always lead to a collision. First, we formally define the notion of STLC that will be of interest to this paper. Let $q(t)$ be the combined state of the robot swarm, where all robot positions are first stacked together in a $2n$ vector, followed by a $n$ vector of robot orientations. Let $p(t)=q(t)[1:2n]$ be the position states of the robots.
@@
The system is small-time locally controllable (STLC) in position states about $p(0)=p_{0},\forall p_{0}\in C_{free}$, if $p(0)=p_{0}$ is contained in the interior of the reachable sets $\mathcal{R}_{p_{0}}$ for all $t>0$. [12]
In Eq. (1), we assumed that the MicroStressBot $i$ has the turning radius ${r_{i}}$. If each robot has a different turning radius, only two groups are needed to achieve STLC through group-based control: group $G_{1}$ that translates all the robots and group $G_{2}$ that rotates all the robots. Now consider the three-step primitive $P_{1}=(G_{1}(d),G_{2}(\pi r_{1}),G_{1}(d))$ where $G_{1}(d)$ moves all the robots forward for $d$, $G_{2}(\pi r_{1})$ rotates all the robots so that robot 1 rotates exactly for $\pi$ radians, and then $G_{1}(d)$ again moves all the robots forward for $d$, making robot 1 return to its starting position. Then the sequence $P_{2}=(P_{1},G_{2}((\pi-\Delta\theta_{2})r_{2}),P_{1})$ moves robots 1 and 2 back to their initial position, where $\Delta\theta_{2}$ is the angular displacement of robot 2 due to $P_{1}$. Iteratively, $P_{n-1}$ moves only one robot while all the others return to the starting position. By first rotating the robot so that the resulting movement due to $P_{n-1}$ is in the desired direction, we can clearly individually move any robot in any direction, showing that the system is STLC. Note that a similar argument has been used in [20] to introduce TSC.
Assuming that all the robots have different turning rates again shifts the burden of control to fabrication and in general scales poorly. In the rest of the paper, we thus assume that all the robots have the same turning rate $r$. We will show that group-control still guarantees STLC.
@@
Given a collection of sets of robots $\{S_{0},S_{1},\ldots\}$ with each $S_{i}$ containing robots with the same turning rate, and $S_{i}$ and $S_{j}$ having different rates if $i\neq j$, STLC can be achieved by combining the approach above with group-control that will be described below within each set.
@@ 4.1 Unilateral Control to Bilateral Control
As stated above, in the rest of the paper, we only consider the case when all the robots have the same turning radius $r$. In [17], we showed that a swarm of MicroStressBots under group-control is (globally) controllable (even without using the group where all the robots rotate). However, we were not able to demonstrate that the system is STLC. In this paper we show that by adding the group where all the robots rotate, STLC can be achieved.
One of the challenges to show that our system is STLC is that the robots are controlled unilaterally. In other words, they can only move forward but not backward. Similarly, they can rotate clockwise but not counter-clockwise. This is the result of the fact that the control input $u$ in Eq. (4) is positive. We show that this restriction can be relaxed.
@@
be the state of the swarm, where $q(t)\in\mathbb{R}^{3n}$. We can see that $q(t)[1:2n]$ are the position states, with $q(t)[2j-1:2j]$ representing the position of robot $j$. Also, $q(t)[2n+1:3n]$ are the orientation states, where $x(t)[2n+j]$ is the orientation of the robot $j$. Next, let $\alpha_{i}=[\alpha_{i,1},\ldots,\alpha_{i,n}]^{T},i=1,\ldots,m$ be the activation vector corresponding to group $i$ being activated. In other words, $\alpha_{i,j}=1$ if robot $j$ belongs to group $i$, and 00 otherwise. The overall swarm dynamics can then be written as:
@@
We introduce the notion of induced vector fields $g_{i}$ and $h_{i}$ so that $f_{i}=g_{i}+h_{i}$. Here, $g_{i}$ contains only the position states information and $h_{i}$ contains only the orientation states information. Given that $G_{m}$ is the group where all the robots rotate, and the rotation radius of all of them is equal, let $f_{m}(\pi)$ represent the rotation of all robots by $\pi$ radians using the rotation-only vector field.
@@
Observe that $f_{i}(d)$ translates the robots that belong to $G_{i}$ for $d$ without changing their orientation, and rotates the robots that do not belong to $G_{i}$ for $\frac{d}{r}$. The application of $f_{m}(\pi)$ rotates all the robots for $\pi$, which means that the robots that belong to the group $G_{i}$ now point in the opposite direction. The second application of $f_{i}(d)$ then brings the robots that belong to $G_{i}$ back to their initial position, while adding another $\frac{d}{r}$ to the orientation of the robots that do not belong to $G_{i}$. Finally, $f_{m}(\pi)$ rotate the robots in $G_{i}$ to their original orientation. The robots not in $G_{i}$ have instead rotated in total for $\frac{2d}{r}+2\pi=\frac{2d}{r}$. ∎
@@
Note that while $h_{i}$ rotates the robots that are not in $G_{i}$ in place, the other robots travel back and forth and could potentially collide with an obstacle. However, we can restrict the motion of the robots in $G_{i}$ to a small ball of the radius $\epsilon$ and apply $h_{i}(\frac{\epsilon}{r})$ multiple times to obtain the desired rotation of the robots not in $G_{i}$ without the robots in $G_{i}$ leaving a small neighborhood of their initial location.
@@
=$\mathcal{F}=\{g_{1},\ldots,g_{m-1},h_{1},\ldots,h_{m-1}\}.$ (6)

### v2

@@ Abstract
This paper investigates how a novel paradigm called group-control can be effectively used for motion planning for microrobot swarms in a global field. We prove that Small-Time Local Controllability (STLC) in robot positions is achievable through group-control, with the minimum number of groups required for STLC being $\log_{2}(n+2)+1$ for $n$ robots. We then discuss the trade-offs between control complexity, motion planning complexity, and motion efficiency. We show how motion planning can be simplified if appropriate primitives can be achieved through more complex control actions. We identify motion planning problems that balance the number of motion primitives with planning complexity. Various instantiations of these motion planning problems are explored, with simulations used to demonstrate the effectiveness of group-control.
@@
Micro Robots Group Control Global Field Control and Motion Planning.
@@ 1 Introduction
Microrobots have gained significant attention for their potential in medical, environmental, and industrial applications. Effective control mechanisms are crucial to enable their diverse functionalities. Various control modalities have been demonstrated for microrobots, including magnetic control [22, 23, 16], optical control [9, 2, 15, 19], and acoustic control [10, 1]. These systems are controlled by global fields, and as a result, they are massively underactuated.
@@
Currently, parallel control of multiple microrobots in a global field is based on the design of robots with distinct physical characteristics, resulting in different responses of each robot to the same control signal. The Global Control Selective Response (GCSR) paradigm [7, 8] uses fabrication to make each robot respond appropriately to the control signal. Turning-rate Selective Control (TSC) [20] is a variation of GCSR and differentiates robots by explicitly designing them with different turning rates. These methods scale poorly to larger swarms due to fabrication complexities. Ensemble Control (EC) [3] leverages differences in the linear velocity and turning rate parameters among robots stemming from randomness in fabrication to control the position of individual robots. Unfortunately, the ability to control individual robots is inherently limited by noise, so this approach also scales poorly.
@@
This paper focuses on electrostatic stress-engineered MEMS microrobots (MicroStressBot), originally described in [6]. The swarm is powered by a uniform electrostatic field generated by a substrate on which the robots move, which means that all robots are controlled by a single global signal. Their size limits onboard control logic and power storage, making it essential to simultaneously coordinate large groups of microrobots for applications like micro-assembly or drug delivery. Our work explores the concept of onboard multi-stage Physical Finite-State Machines (PFSMs) introduced in [21]. PFSMs allow robots to be individually addressed and activated one by one. Our previous work [17] takes the idea of PFSMs further by introducing group-based control, where several robots can be activated together as a group. This paper investigates the trade-off between the motion planning complexity, control complexity, and motion efficiency within the group-control framework. In particular, we introduce several control and motion planning problems and discuss their scalability.
@@
Motion planning for microrobots involves determining a sequence of movements to reach a given configuration while avoiding obstacles and collisions. Various methods exist, including graph-based methods like Dijkstra’s and $A^{*}$ search [13], sampling-based algorithms like Rapidly-exploring Random Trees (RRTs) [14] and optimization-based methods [11, 29]. Recent advances in machine learning, such as [24, 26, 27], leverage neural networks to speed up RRTs and control techniques for mobile robot navigation. However, despite a wealth of solutions for motion planning problems, motion planning for microrobot swarms in a global field remains a challenging problem because the system is high-dimensional yet massively underactuated, having a single control signal.
@@
Fundamentally, when a global control signal controls a microrobot swarm, the robots can be controlled individually only when each robot responds to the control signal differently. In this paper, we introduce the group-control framework to differentiate the motion of each microrobot in the swarm. We prove that Small-Time Local Controllability (STLC) is achievable in robot positions given an appropriate group allocation, with the minimum number of groups required for STLC being $log_{2}(n+2)+1$ for $n$ robots. Additionally, we discuss the complexity trade-offs between control, motion planning, and motion efficiency. We identify motion planning problems that balance the number of robot groups and motion primitives with planning complexity. Various instantiations of these motion planning problems are explored, and simulations are provided that demonstrate the scalability and effectiveness of the proposed algorithms.
@@ 2.1 MicroStressBot
The MicroStressBot is a 120 $\mu$m $\times$ 60 $\mu$m $\times$ 10 $\mu$m mobile microrobot platform introduced in [6]. A MicroStressBot has two actuated internal degrees of freedom (DOF): an untethered scratch-drive actuator (USDA) that provides forward motion and a steering-arm actuator that determines whether the robot moves in a straight line or rotates (Fig. 1). A single MicroStressBot can have its arm raised or lowered, depending on the voltage applied across a substrate formed by an electrode array. When a sufficiently high voltage is applied to the substrate, the arm is pulled into contact with the substrate, and the robot rotates around the contact point. In contrast, when the voltage is reduced below a threshold, the arm is raised, and the robot moves forward.
MicroStressBot control has been successfully implemented in [8]. It has been shown that if the pull-down and release voltages of the robots are different, they can be independently controlled. However, it is difficult to consistently fabricate robots to respond in the desired way. As a result, the approach scales poorly.
Figure 1: The schematic of a MicroStressBot.
@@ 2.2 Physical FSM Robots
One solution to dramatically improve the scalability of the microrobot swarm control is to make the robots respond to a temporal sequence of (a small number of) voltage levels rather than to the voltage directly. Finite-State Machines (FSMs) can accept a set of input sequences [25] (sequences of control signal levels). Previously, in [21], we proposed the on-board MEMS Physical FSM (PFSM) that, upon acceptance of a unique control signal sequence, causes the arm of the MicroStressBot to be pulled up or down. PFSMs can be constructed from several basic modules that are combined together and thus fabricated efficiently. In this work, we build on this idea to propose group-control. In particular, we use the fact that several PFSM modules, each corresponding to one group, can be combined together so that each robot can belong to several groups.
@@ 2.3 Group-Control
The core idea of group-control is that by equipping the robots with PFSMs, they can be assigned to (several) different groups. When the activation sequence corresponding to a group is sent to the swarm, all the robots that belong to the group are activated. In this paper, activating a MicroStressBot corresponds to raising its arm so that it can move forward (translate); otherwise, the robot rotates. By assigning each robot to a unique subset of groups, we can differentiate between the robots and make them respond differently to a sequence of inputs that activate all the groups one after the other, each time moving the robots belonging to the group for some distance.
Table 1 shows how six robots can be assigned to different groups. To make the selection of groups unique to each robot, if we have $n$ robots, we need $m=\log_{2}(n+2)+1$ groups ($n+2$ rather than $n$ because each robot needs to belong to at least one group so it can move forward, and no robot can belong to all the groups so it can rotate). We add one more group where none of the robots translate, they all rotate in place. This special group will be the group $G_{m}$.
As can be easily seen from Table 1, group assignment corresponds to assigning a unique bit pattern (a subset of groups) to each robot, where each bit corresponds to a group (labeled $G_{i}$, $i=1,2,3$ in Table 1). The robot then belongs to the groups where the corresponding bit equals $1$. For example, robot $R_{1}$ belongs to the group $G_{3}$, robot $R_{3}$ belongs to the groups $G_{2}$ and $G_{3}$, etc. The group $G_{4}$ is the additional group with all zeros where all the robots are rotating. It is clear that this method guarantees that each robot has a distinct group allocation. Also, note that when $n=2^{m-1}-2$ (the maximum number of robots that can be controlled by $m$ groups), each group (apart from the special group $G_{m}$) contains exactly half, that is, $\frac{n}{2}=2^{m-2}-1$, robots.
@@
The group allocation can be realized by the on-board PFSMs. At each time step $k$, only one group of robots is activated, and the robots belonging to that group translate (move forward) while the remaining robots rotate. We call a sequence of group selections an activation sequence. If $m$ is the number of groups and $n$ is the number of robots, we would need $k=O(\log_{2}m)=O(\log_{2}(\log_{2}(n+2)+1))$ bits to select a group through a PFSM [21]. Thus, PFSMs for group-based control can be significantly simpler than when each robot needs to be selected individually.
@@ 3.1 Dynamics
To describe group-control mathematically, we start with a dynamic model of a MicroStressBot. The robot can move freely on a horizontal plane, so its configuration space is the Euclidean group $SE(2)$. We will describe the state of the robot $i$ with a vector $[x_{i},y_{i},\theta_{i}]^{T}$, where $x_{i}$ and $y_{i}$ are the Cartesian coordinates of the pivot point of the robot and $\theta_{i}$ is the robot orientation. The equations of motion for the robot $i$ are given by:
$\frac{d}{dt}\begin{bmatrix}x_{i}\\ y_{i}\\ \theta_{i}\end{bmatrix}=a_{i}\cdot\begin{bmatrix}\cos\theta_{i}\\ \sin\theta_{i}\\ 0\end{bmatrix}\cdot u+(1-a_{i})\cdot\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}\cdot\frac{u}{r_{i}}$ (1)
where $a_{i}\in\{0,1\}$ is the switching input that determines whether the robot $i$ is moving forward or turning in place, $u\in\mathbb{R}^{+}$ is the rate of forward motion, and $r_{i}$ is the turning radius of the robot $i$. If the control input is piecewise constant over each epoch $\Delta T$, we obtain the following discrete-time model:
@@
Note that $u$ is a unilateral control input. In other words, MicroStressBot cannot go backward or turn clockwise.
@@ 3.2 Switched and Embedded Systems
We introduce new variables $v_{i}(t)\in[0,1]$ that satisfy $\sum_{i=1}^{m}v_{i}(t)=1$. This transforms the switched system (3) to the embedded form [4, 28]:
@@
where $u_{i}$ is the control input for each vector field $f_{i}$. It can be shown [4] that the set of trajectories of the switched system (3) is a dense subset of the set of trajectories of the embedded system (4). Consequently, any trajectory of the embedded system can be approximated by the switched system to any desired accuracy, which means that the controllability of the switched system is equivalent to the controllability of the embedded system.
@@ 4 Controllability Analysis
For the robots to navigate among obstacles, the system, in general, needs to be Small-Time Locally Controllable (STLC). Otherwise, there are configurations in the obstacle-free configuration space $C_{free}$ that will always lead to a collision. First, we formally define the notion of STLC that will be of interest in this paper. Let $q(t)$ be the combined state of the robot swarm, where all the robot positions are first stacked together in a $2n$ vector, followed by a $n$ vector of the robot orientations. Let $p(t)=q(t)[1:2n]$ be the position states of the robots, and let $\Theta(t)=q(t)[2n+1:3n]$ be the orientations. Let $\mathcal{R}_{(p_{0},\Theta_{0})}(T)$ be the set of all positions reachable by the trajectories of the system starting at $q(0)=(p_{0},\Theta_{0})$ in time no greater than $T$.
@@
The system is small-time locally controllable (STLC) in position states about an initial state $(p_{0},\Theta_{0})$, if $p_{0}$ is contained in the interior of the reachable sets $\mathcal{R}_{(p_{0},\Theta_{0})}(T)$ for all $T>0$. [12]
Note that each MicroStressBot in the swarm has $SE(2)$ symmetry (if the initial configuration of the robot is moved, its trajectories will move by the same amount). So, STLC at one initial configuration implies STLC everywhere.
In Eq. (1), we assumed that MicroStressBot $i$ has the turning radius ${r_{i}}$. If each robot has a different turning radius, only two groups are needed to achieve STLC through group-based control: group $G_{1}$ that translates all robots and group $G_{2}$ that rotates all robots. To see this, consider the three-step primitive $P_{1}=(G_{1}(d),G_{2}(\pi^{r_{1}}),G_{1}(d))$ where $G_{1}(d)$ moves all robots forward for $d$, $G_{2}(\pi^{r_{1}})$ rotates all robots so that robot 1 rotates exactly for $\pi$ radians, and then $G_{1}(d)$ again moves all the robots forward for $d$, making robot 1 return to its starting position. Then the sequence $P_{2}=(P_{1},G_{2}((\pi-\Delta\theta_{2})^{r_{2}}),P_{1})$ moves robots 1 and 2 back to their initial position, where $\Delta\theta_{2}$ is the angular displacement of robot 2 due to $P_{1}$. Iteratively, $P_{n-1}$ moves only one robot while all others return to the starting position. By first rotating the robot so that the resulting movement due to $P_{n-1}$ is in the desired direction, we can individually move any robot in any direction, showing that the system is STLC. Note that a similar argument has been used in [20] to introduce TSC.
Assuming that all the robots have different turning rates again shifts the burden of control to fabrication and generally scales poorly. In the rest of the paper, we thus assume that all robots have the same turning rate $r$. We will show that group-control still guarantees STLC.
@@
Given a collection of sets of robots $\{S_{0},S_{1},\ldots\}$ with each $S_{i}$ containing robots with the same turning rate, and $S_{i}$ and $S_{j}$ having different rates if $i\neq j$, STLC can be achieved by combining the approach above with the methodology that will be described below within each set.
@@ 4.1 Unilateral Control to Bilateral Control
As stated above, in the rest of the paper, we only consider the case where all the robots have the same turning radius $r$. In [17], we showed that a swarm of MicroStressBots under group-control is (globally) controllable (even without using the group where all the robots rotate). However, we were unable to demonstrate that the system is STLC. In this paper, we show that by adding the group where all the robots rotate, STLC can be achieved.
One of the challenges in showing that our system is STLC is that the robots are controlled unilaterally. In other words, they can only move forward but not backward. Similarly, they can rotate counter-clockwise but not clockwise. This is the result of the control input $u$ in Eq. (4) being positive. We show that this restriction can be relaxed.
@@
be the state of the swarm, where $q(t)\in\mathbb{R}^{3n}$. We can see that $q(t)[1:2n]$ are the position states, with $q(t)[2j-1:2j]$ representing the position of robot $j$. Also, $q(t)[2n+1:3n]$ are the orientation states, where $x(t)[2n+j]$ is the orientation of the robot $j$. Next, let $\alpha_{i}=[\alpha_{i,1},\ldots,\alpha_{i,n}]^{T},i=1,\ldots,m$ be the activation vector corresponding to the group $i$ being activated. In other words, $\alpha_{i,j}=1$ if robot $j$ belongs to group $i$, and 00 otherwise. The overall swarm dynamics can then be written as follows:
@@
We introduce the notion of induced vector fields $g_{i}$ and $h_{i}$ such that $f_{i}=g_{i}+h_{i}$. Here, $g_{i}$ contains only position state information ($f_{i}[2n+1:3n]=0$), and $h_{i}$ contains only the orientation state information ($h_{i}[1:2n]=0$). Given that $G_{m}$ is the group where all the robots rotate and their rotation radius is the same, let $f_{m}(\pi)$ represent the rotation of all the robots by $\pi$ radians using the rotation-only vector field.
@@
Observe that $f_{i}(d)$ translates the robots that are in $G_{i}$ for $d$ without changing their orientation and rotates the robots that are not in $G_{i}$ for $\frac{d}{r}$. The application of $f_{m}(\pi)$ rotates all the robots for $\pi$, which means that the robots that are in $G_{i}$ now point in the opposite direction. The second application of $f_{i}(d)$ then returns the robots that are in $G_{i}$ to their initial position while adding another $\frac{d}{r}$ to the orientation of the robots that are not in $G_{i}$. Finally, $f_{m}(\pi)$ rotate the robots in $G_{i}$ to their original orientation. The robots not in $G_{i}$ have instead rotated in total for $\frac{2d}{r}+2\pi=\frac{2d}{r}$. ∎
@@
Note that while $h_{i}$ rotates the robots that are not in $G_{i}$ in place, the other robots travel back and forth and could potentially collide with an obstacle. However, we can restrict the motion of the robots in $G_{i}$ to a small ball of radius $\epsilon$ and apply $h_{i}(\frac{\epsilon}{r})$ multiple times to obtain the desired rotation of the robots not in $G_{i}$ without the robots in $G_{i}$ leaving a small neighborhood of their initial location.
@@
=$\mathcal{F}=\{g_{1},\ldots,g_{m-1},h_{1},\ldots,h_{m-1},f_{m}\}.$ (6)